OpenStudy (anonymous):

Evaluate lim-->2 sqaure root of x -sqaure root of 2, divided by 3x-6?

myininaya (myininaya):

\[\lim_{x \rightarrow 2}\frac{\sqrt{x}-\sqrt{2}}{3x-6}=\lim_{x \rightarrow 2}\frac{\sqrt{x}-\sqrt{2}}{3(x-2)}=\lim_{x \rightarrow 2}\frac{\sqrt{x}-\sqrt{2}}{3(\sqrt{x}-\sqrt{2})(\sqrt{x}+\sqrt{2})}\]

myininaya (myininaya):

\[\lim_{x \rightarrow 2}\frac{1}{3(\sqrt{x}+\sqrt{2})}\]

myininaya (myininaya):

not you can plug in 2